You mix 125 mL of 0.270 M CsOH with 50.0 mL of 0.675 M HF in a coffee-cup calorimeter, and the temperature of both solutions rises from 20.70 °C before mixing to 23.82 °C after the reaction.

CsOH (aq) + HF (aq) ---> CsF (aq) + H2O (l)

What is the enthalpy of reaction per mole of CsOH ? Assume the densities of the solutions are all 1.00 g/mL, and the specific heat capacities of the solutions are 4.2 J/g · K.

Enthalpy of reaction =  kJ/mol
Answer #1

Total volume = 125 + 50 = 175 ml

Mass = volume * density

= 175 *1 = 175 grams

Q = mass * specific heat * ∆T

= 175 * 4.2 * ( 23.82 - 20.70 )

= 2293.2 joules

Now moles of CsOH = molarity * volume ( in litre )

= 0.27 * 0.125 ( in litre )

= 0.03375 moles

Now 0.03375 mole produces 2293.2 joules

1 mole will produce = 2293.2 / 0.03375

= 67946.67 joules

In KJ / mole = 67.95 KJ / mole

If you have any query please comment

If you satisfied with the solution please like it thankxx